OpenStudy (anonymous):

Proof that any multiple of a thousand greater or equal than 5000 can be expressed as the sum of some multiple of 5000 and some multiple of 2000 So, the proof is supposed to be by induction. I've checked the base case, 5000= 5000*1+2000*0 I got the hipotesis 1000k = 5000a+2000b But i can't figure out how to use that in order to prove the proposition for k+1. Thanks.

OpenStudy (perl):

we can prove this

OpenStudy (perl):

assume the statement is true for n=k. so 1000k = 5000*a + 2000*b , for k>=5

OpenStudy (perl):

then prove for n=k+1

OpenStudy (perl):